OpenStudy (anonymous):

Part A: The area of a square is (25x2 - 20x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. Part B: The area of a rectangle is (49x2 - 36y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. Part C: The volume of a rectangular box is (x3 + 2x2 - 9x - 18) cubic units. Determine the dimensions of the rectangular box by factoring the volume expression completely. Show your work.

OpenStudy (anonymous):

@Gebooors

OpenStudy (gebooors):

Try (5x-2)^2

OpenStudy (anonymous):

what part is that for? try that for what? why ?

OpenStudy (gebooors):

That is for A

OpenStudy (anonymous):

can you elaborate

OpenStudy (gebooors):

Side of square is 5x -2, sides are equal. Try to calculate (5x -2)^2 You can use binomical expansion

OpenStudy (anonymous):

i really dont and cannot

OpenStudy (anonymous):

understand at least

OpenStudy (gebooors):

Are you familiar with (a+b)^2 = a^2 + 2ab + b^2

OpenStudy (gebooors):

Then (a -b) ^2 = a^2 -2ab + b^2

OpenStudy (anonymous):

no /:

OpenStudy (gebooors):

And third one (a +b) (a - b)= a^2 -b^2

OpenStudy (gebooors):

I'll show you:

OpenStudy (gebooors):

(5x -2)^2 =(5x -2)(5x -2) = 5x * 5x + 5x *(-2) + (-2) * 5x+ (-2)(-2) Can you continue from here?

OpenStudy (anonymous):

not really

OpenStudy (gebooors):

You start to Multiply: 5x * 5x = 25*x^2

OpenStudy (anonymous):

25x=25x2

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Are you going to work on it, this time?

OpenStudy (tkhunny):

(25x2 - 20x + 4) This should be recognized as a perfect square trinomial. As a perfect square, the pieces had better be the square roots of the terms we already have. 25x^2 ==> 5x 4 ==> 2 Thus (5x + 2)^2 and 5x+2 is the length of the side of the square.

OpenStudy (anonymous):

thats it for part a ? @tkhunny

OpenStudy (anonymous):

and what about part b

OpenStudy (tkhunny):

For part B, you must recognize the Difference of Squares and factor it as such. It wirks like this. \((a^{2} - b^{2}) = (a+b)(a-b)\) Go!

OpenStudy (anonymous):

lol good one

OpenStudy (tkhunny):

You may wish to consider factoring by grouping. x3 + 2x2 - 9x - 18 x^2(x+2) - 9(x+2) = (x^2 - 9)(x+2) Finally, your task is to recognize the difference of squares and continue with the factoring - just like the last problem.

OpenStudy (tkhunny):

(7x-6y)(7x+6y) -- That is correct.

OpenStudy (tkhunny):

I regret that I get only one Medal per thread. EXCELLENT work. Learn a little more mathematics and you will be able to tell when your accountant and money manager are trying to rip you off and blow all your hard-earned millions.

OpenStudy (anonymous):

i only know the answers cuz i got them from someone else lol however i know absolutely nothing about this very fun very useful math

OpenStudy (tkhunny):

...and I thought we were getting along a little better, this time. {wry smile}{tiny chuckle}

OpenStudy (anonymous):

oh no, never with math in the way

OpenStudy (the_fizicx99):

Tk, notice how 25x^2 - 20x + 4 has a minus sign, √25x^2 = 5x √4 = 2 (5x - 2)^2 = (5x)^2 -2(5x)(2) - (2)^2 = 25x^2 -20x + 4 \(\ \color{green}\checkmark\)

OpenStudy (tkhunny):

Whoops! So it does. Good call.